Lina had a piece of tape 4 1/2 meters long. She used 2/5 of it. What part of the tape was not used?

Answer by Theo(13342)

the tape was 4.5 meters long.

she used 2/5 of it = 2/5 * 4.5 = 1.8 meters.

the part of the tape not used was 3/5 of it = 2.7 meters.

1.8 + 2.7 = 4.5, confirming the solution is correct.

the solution is that the part of the tape not used is 2.7 meters.